OpenStudy (anonymous):

Prove that the rate of heat production in each of the two resistors connected in parallel are inversely proportional to the rsistances.

OpenStudy (anonymous):

Hai sonia...if i understood ur ques it seems simple the rate of heat production in any conductor can be proved to be \[P=V \times I\] same as \[V ^{2}\div R\] (ill give the proof if u need) then for a parallel combination of resistors \[P1=V^2/R1\] & \[P2=V^2 /R2\] The inverse relationship in each of the resistors to to the resistance is obvious from the term R2 appearing in the denominator